# Automated Data Augmentation for Few-Shot Time Series Forecasting: A Reinforcement Learning Approach Guided by a Model Zoo

### v2

@@ 1 Introduction
Time series forecasting is a critical task with diverse applications, but it faces significant challenges due to the limited availability of high-quality training data. This challenge is further amplified in time-evolving domains with inherent non-stationarity and becomes even more pronounced in few-shot learning scenarios, where the scarcity of data severely limits the performance of forecasting models. While recent advancements have focused on developing more effective deep-learning architectures to capture long-term trends, clinical patterns, and multivariate relationships in the data (Wu et al., 2021; Nie et al., 2023; Liu et al., 2024b), in this work, we explore a data-centric approach and introduce a novel, learning-based data augmentation method that can be seamlessly integrated with existing forecasting models.
@@
We recognize that effective data augmentation requires generating high-quality, diverse training samples. However, in practice, existing forecasting models typically rely on fixed-form data augmentation techniques (Wen et al., 2021; Cheung & Yeung, 2020), which lack data-dependent adaptability and may introduce unexpected noises. Previous learning-based augmentation methods typically involve contrastive learning (Demirel & Holz, 2024) or Mixup combinations (Schneider et al., 2024) to generate new data sequences. However, these methods are not task-specific, as the data generation process is not guided by the performance of forecasting models. In contrast, we argue that aligning the training objectives of augmentation models with the resulting forecasting performance is essential, enabling a closed-loop optimization of the data augmentation process.
@@
In this work, we present ReAugment, which uses reinforcement learning (RL) to create the closed-loop data augmentation process. Our goal is to address data scarcity and enhance the generalizability of forecasting models in few-shot learning scenarios. We aim to answer three critical questions: which parts of the training set should be augmented, how the augmentation should be performed, and what advantages RL offers compared to previous approaches. Specifically, ReAugment dynamically identifies a training subset of overfit-prone data samples that would benefit most from augmentation and automatically searches for optimal augmentation policies tailored to these samples.
@@
ReAugment initially leverages a forecasting model zoo, consisting of diverse instances of the same network architecture trained via cross-validation, to identify overfit-prone training samples in need of augmentation. An interesting finding is that training forecasting models exclusively with data samples that present low prediction variance across the model zoo often leads to better performance (detailed in Section 3). This insight leads us to select the data points with high prediction variance across the model zoo as the root of the subsequent augmentation process. Additionally, targeting overfit-prone samples for augmentation ensures more efficient use of limited data resources, compared to applying uniform augmentation to all training samples.
@@
Based on the identified overfit-prone samples, ReAugment learns an augmentation network in forms of a variational masked autoencoder (VMAE), where we use the latent space as the action space to be optimized in the RL framework. The RL-based augmentation policies are guided by a reward function derived from the backtesting errors, measured by applying the generated data to the forecasting model zoo. The training objective balances both the diversity and quality of the augmented data: It encourages the new data points to fall into regions where the forecasting models tend to overfit, while maintaining alignment with the original data distribution. Specifically, we employ the REINFORCE algorithm (Williams, 1992) to enable the backtesting diversity evaluated on the model zoo, which is non-differentiable, to guide the learning process of the augmentation network.
The contributions of this work are as follows:
Key idea: We propose ReAugment, which offers a pilot study on leveraging RL for time series augmentation.
Finding anchor data: We propose a novel method for identifying overfit-prone samples by leveraging cross-validation errors from a forecasting model zoo to pinpoint anchor points with potential value for data augmentation.
Model: We introduce a novel network architecture based on a variational masked autoencoder for sequential data, where the latent space is used as the action space of RL.
Algorithm: We define a specific reward function that incorporates the model zoo’s prediction responses to guide the generation of new data points targeting regions where the forecasting models are prone to overfitting.
@@
Results: ReAugment demonstrates substantial improvements across a wide range of base models and datasets, especially in few-shot learning scenarios.
@@ 2 Related Work
Compared with CNN-based or RNN-based forecasting models (Torres et al., 2021; Wang et al., 2022; Che et al., 2018; Sagheer & Kotb, 2019), recent Transformer-based methods (Li et al., 2019; Wu et al., 2021; Zhou et al., 2021; Liu et al., 2021; Zhou et al., 2022; Zhang & Yan, 2022; Zeng et al., 2023; Cao et al., 2024; Yi et al., 2024) have shown superior performance across a wide range of time series forecasting tasks. Most previous work has focused on developing more effective network architectures to capture long-term trends, clinical patterns, and multivariate relationships within the data. For instance, Autoformer (Wu et al., 2021) improves upon the Transformer by employing a deep decomposition architecture that progressively separates trend and seasonal components throughout the forecasting process. PatchTST (Nie et al., 2023) vectorizes time series data into patches of specified size, which are then encoded through a Transformer, with the model producing forecasts of the desired length via an appropriate prediction head. LSTF-Linear (Zeng et al., 2023) simplifies complex time series forecasting problems and outperforms many Transformers by using a set of remarkably simple one-layer linear models. iTransformer (Liu et al., 2024b) modifies the architecture by adopting components with inverted dimensions, demonstrating superior performance on multivariate time series data. Unlike these methods, we explore a data-centric approach to time series forecasting, rather than focusing on the model architecture. We introduce a data augmentation method that can be seamlessly integrated with existing forecasting models.
@@
Few-shot time series forecasting refers to the ability to make accurate predictions for time series data with very limited historical information. Traditional time series forecasting models, such as ARIMA, Exponential Smoothing, and deep learning models like LSTMs or the above Transformer-based models, generally require large amounts of historical data to make reliable predictions. However, in many real-world applications, obtaining sufficient data can be challenging, especially in scenarios where data is scarce, noisy, or difficult to collect (Dooley et al., 2023; Xu et al., 2024; Jiang et al., 2023; Yuan et al., 2024). Notably, recent literature has introduced large foundation models specifically designed for time series forecasting (Das et al., 2024; Jin et al., 2024; Bian et al., 2024; Ekambaram et al., 2024; Liu et al., 2024a, c; Pan et al., 2024), often evaluating these models under zero-shot domain generalization settings. However, in our preliminary experiments, we found that as the distribution gap between training and testing data increases (e.g., when data comes from entirely different domains), the generalization performance of these models significantly decreases.
@@
Iglesias et al. (2023) have presented a taxonomy of augmentation techniques. Simple augmentation methods involve time, frequency, and magnitude domain transformation techniques such as slicing (Cao et al., 2020), frequency warping (Cui et al., 2015), and jittering (Flores et al., 2021). The second category is the learning-based methods, including those based on contrastive learning (Demirel & Holz, 2024), and Mixup combinations (Schneider et al., 2024). Additionally, advanced generative models have been employed to generate realistic time series data, including the GAN-based (Yoon et al., 2019; Liao et al., 2020), VAE-based (Sohn et al., 2015; Li et al., 2020), and Diffusion-based (Huang et al., 2023) methods. The generated samples can be used for further training of the forecasting models. In contrast to existing approaches, we present a pilot study on using reinforcement learning for time series data augmentation.
@@ 3 Overfit-Prone Samples as Anchor Data
The first challenge we need to address is identifying which parts of the training set would benefit the most, so they can be used as anchor points in the augmentation process.
@@ 3.1 Finding Overfit-Prone Data with a Model Zoo
A key assumption in time series forecasting, particularly in few-shot learning scenarios, is that forecasting models tend to overfit certain regions of the training data. Different training samples can have varying impacts on training quality. Intuitively, we aim to generate new data with distributions around the original data points that have higher probabilities of overfitting. To achieve this, we measure the cross-validation errors from a batch of forecasting models to pinpoint the overfit-prone samples.
To construct the forecasting model zoo, we divide the training set into parts and perform -fold cross-validation. In this way, we obtain sets of model parameters for the same network architecture, e.g., iTransformer (Liu et al., 2024b), which we denote as . For each data point , we evaluate the prediction errors of the other models that were not trained on this subset, along with the training error of the model trained on this subset. We then calculate the variance of the MSE across the models, denoted as , and sort the “model-zoo variance” of all data points.
The model-zoo variance measures the inconsistency in predictions made by different models trained on slightly different subsets, and can therefore serve as an indicator of the overfit-prone samples. When a model overfits, its predictions on certain data points are highly sensitive to the specific subset it was trained on, leading to a high variability of the performance across the models. As a result, data points with high model-zoo variance are more likely to be in regions where the models struggle to generalize, making them overfit-prone and thus ideal candidates for data augmentation to improve generalization.
In practice, we split the training set into two subsets based on the top and bottom values of the model-zoo variance, and refer to the top subset as the overfit-prone samples.
@@ 3.2 Preliminary Findings on Overfit-Prone Data
To understand the impact of the overfit-prone samples on the training process, we conduct the following experiments. We train forecasting models separately using data from Group A (the subset with top model-zoo variance) and Group B (the subset with bottom model-zoo variance). We evaluate the models on the same test set. Given its superior average performance, we select iTransformer as the preferred model for these experiments. The experiments are conducted on classic real-world multivariate time series benchmarks, including ETT, Weather, and Electricity.
The forecasting results are shown in Figure 2, where a lower MAE indicates more accurate predictions. Our experiments demonstrate that the choice of training sets has a significant impact on final performance. Predictions trained in Group B consistently outperform those trained in Group A across all benchmarks, with substantial margins. This finding suggests that the overfit-prone samples are more likely to negatively affect the training quality of forecasting models. Since data augmentation is an effective strategy to mitigate overfitting, we propose generating new data around the distributions of the overfit-prone samples, allowing the forecasting model to learn more generalizable patterns from these data points.
@@ 4.2 Variational Masked Autoencoder as the RL Actor
We design a Variational Masked Autoencoder (VMAE), as illustrated in Figure 3(left), which takes masked time series data with absolute timestamp in the whole dataset as input and outputs the complete corresponding data. The entire architecture contains four modules, including (i) the prior network, which learns the prior distribution of based on the masked data, (ii) the posterior module, which learns the posterior distribution of based on the original data, (iii) the data encoder, which extracts significant features from with , (iv) the decoder, which generates data from the encoding features and the latent variables. These modules are parametrized by and . We have
| Prior: | (1) | |||
| Posterior: | ||||
| Encoder: | ||||
| Decoder: |
@@
In general, the VMAE design can be integrated into any encoder-decoder-based time series forecasting architecture. In this work, we specifically adopt the encoder and decoder (implemented as a linear projector) from iTransformer (Liu et al., 2024b) for feature extraction and decoding.
4.3 REINFORCE Guided by Model Zoo Predictions
The fundamental idea behind training Stage B is that simply generating data by approximating the distribution of the original data is insufficient. Therefore, we introduce an RL framework to expand the data distribution of overfit-prone samples, with the hope that it will naturally cover important patterns in the test set and prevent the forecasting from learning trivial solutions to fit the original few-shot data.
Intuitively, we want the augmented data to have two properties: (i) it should be reasonably diversified to increase the data size and broaden the distribution around overfit-prone samples, and (ii) it should not deviate too far from the original distribution, as excessive divergence may introduce undesired noise. To achieve the first goal, inspired by our preliminary findings, a straightforward approach is to augment the dataset with samples that exhibit diverse prediction results across the model zoo. The insight is to augment the overfit-prone data with the potential overfit-prone data. The model zoo, denoted by , consists of instances of the same network architecture, obtained through cross-validation with pretrained parameters .
However, a practical challenge is that the “model-zoo variance” is not differentiable and thus cannot be directly used to optimize the augmentor through gradient descent. To tackle this problem, we propose to finetune the neural augmentor using a REINFORCE algorithm, taking the latent space generated by the prior module in VMAE as the action space. As illustrated in Figure 3, we incorporate the model-zoo variance evaluated on the augmented data as part of the reward function:
where represents the prediction results from the -th pretrained instance. Subsequently, we finetune the VMAE’s prior module through REINFORCE, treating the prior learner as the policy network while keeping the other network parameters fixed. The learning objective is to maximize the following reward function:
The reward function encourages augmented samples to exhibit more pronounced predictive variance across the model zoo, while also constraining the differences between the augmented and original data. In the reward function, a scaled-sigmoid function is employed to minimize the likelihood of rewards clustering around or controlled by hyperparameter . This approach ensures that, despite potential order-of-magnitude differences in reconstruction error and backtest variance within the model zoo, the reward function can learn effectively from these variations.
Based on this reward function, the policy network (i.e., the prior network in VMAE) is optimized as follows via gradient ascent, where is the learning rate:
@@ 5 Experiments
| Training Data | iTransformer (2024b) | PatchTST (2023) | DLinear (2023) | Average | ||||
@@
| Dataset | Original | Gaussian | Convolve | TimeGAN | ADA | ReAugment | Promotion | |||||||
| MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | |
| ETTh1 | 0.434 | 0.411 | 0.437 | 0.416 | 0.441 | 0.417 | 0.444 | 0.419 | 0.435 | 0.413 | 0.4220.01 | 0.4030.01 | 2.76% | 1.95% |
| ETTh2 | 0.362 | 0.320 | 0.365 | 0.321 | 0.364 | 0.323 | 0.366 | 0.327 | 0.368 | 0.331 | 0.3390.01 | 0.3020.01 | 6.35% | 5.63% |
| ETTm1 | 0.440 | 0.470 | 0.438 | 0.469 | 0.426 | 0.479 | 0.430 | 0.483 | 0.429 | 0.484 | 0.4100.01 | 0.4360.01 | 6.82% | 7.23% |
| ETTm2 | 0.282 | 0.204 | 0.283 | 0.204 | 0.286 | 0.207 | 0.285 | 0.206 | 0.284 | 0.207 | 0.2750.02 | 0.1960.01 | 2.48% | 3.92% |
| Weather | 0.231 | 0.187 | 0.240 | 0.196 | 0.253 | 0.204 | 0.239 | 0.191 | 0.246 | 0.198 | 0.2290.00 | 0.1850.00 | 0.87% | 1.07% |
| Electricity | 0.258 | 0.168 | 0.263 | 0.170 | 0.262 | 0.170 | 0.267 | 0.177 | 0.265 | 0.171 | 0.2540.01 | 0.1650.01 | 1.55% | 1.79% |
| Traffic | 0.318 | 0.466 | 0.319 | 0.467 | 0.320 | 0.463 | 0.315 | 0.449 | 0.318 | 0.456 | 0.2930.01 | 0.4290.01 | 7.86% | 7.94% |
| Exchange | 0.228 | 0.103 | 0.229 | 0.104 | 0.226 | 0.099 | 0.226 | 0.100 | 0.235 | 0.116 | 0.2240.00 | 0.0970.00 | 1.75% | 5.83% |
| Dataset | Original | Guassian | Convolve | TimeGAN | ReAugment | |||||
| MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | |
| ETTh1 | 0.405 | 0.387 | 0.4070.02 | 0.3920.01 | 0.416 | 0.399 | 0.409 | 0.390 | 0.3960.01 | 0.3810.01 |
| ETTh2 | 0.350 | 0.301 | 0.3520.01 | 0.3070.01 | 0.356 | 0.303 | 0.348 | 0.299 | 0.3460.01 | 0.2940.01 |
| ETTm1 | 0.377 | 0.341 | 0.3740.02 | 0.3400.02 | 0.387 | 0.352 | 0.392 | 0.357 | 0.3640.01 | 0.3280.01 |
| ETTm2 | 0.272 | 0.186 | 0.2720.01 | 0.1870.00 | 0.275 | 0.188 | 0.279 | 0.190 | 0.2630.01 | 0.1790.00 |
| Weather | 0.219 | 0.178 | 0.2270.01 | 0.1870.01 | 0.265 | 0.210 | 0.219 | 0.177 | 0.2060.00 | 0.1700.00 |
| Electricity | 0.239 | 0.148 | 0.2430.01 | 0.1500.00 | 0.264 | 0.170 | 0.276 | 0.183 | 0.2360.01 | 0.1470.01 |
| Traffic | 0.269 | 0.392 | 0.2690.01 | 0.3940.02 | 0.283 | 0.407 | 0.296 | 0.412 | 0.2640.01 | 0.3880.01 |
| Exchange | 0.206 | 0.086 | 0.2080.00 | 0.0870.00 | 0.210 | 0.087 | 0.210 | 0.087 | 0.2040.00 | 0.0850.00 |
5.1 Experimental Setups
Following previous work (Wu et al., 2021; Liu et al., 2024b; Nie et al., 2023; Zeng et al., 2023), we thoroughly evaluate the proposed ReAugment on five publicly available real-world datasets: ETT (including subsets), Traffic, Electricity, Weather, and Exchange. Please refer to Appendix B for more details on these datasets.
We evaluate ReAugment in two setups: the few-shot learning setup and the standard setup, which provides full access to the original training set of the above datasets.
First, in the few-shot setup, we simulate scenarios with limited training data to assess the model’s ability to handle data scarcity. Specifically, we reduce the training set size to either or of the full dataset, depending on the dataset characteristics. The few-shot training data corresponds to the earliest portion of the time series, ensuring a significant distribution shift from the test set. The validation and test sets remain consistent with those used in prior studies.
Second, in the standard setup, we follow the configuration from previous work, allowing access to the full training set while keeping the validation and test sets unchanged.
We primarily use iTransformer (Liu et al., 2024b) for the forecasting model, as it has demonstrated strong performance in standard time series forecasting tasks. We also conduct experiments with another Transformer-based method, PatchTST (Nie et al., 2023), and the linear model DLinear (Zeng et al., 2023). Unless otherwise specified, we use a model zoo consisting of cross-validation models.
We employ a fixed lookback length of time steps across all datasets and report the multivariate sequence prediction results with prediction lengths of time steps. We provide additional implementation details in Appendix C, including information on hyperparameters, hardware requirements, and computational costs.
5.2 Baseline Augmentation Methods
Gaussian augmentation.
@@ 5.2 Baseline Augmentation Methods
Inspired by prior literature (Iglesias et al., 2023), we use traditional data augmentation methods that apply simple transformations, such as adding Gaussian noise to the raw data. This approach enhances the diversity of the training data by adjusting the controllable variances and means of the added Gaussian noise.
Convolve augmentation.
@@
We employ another traditional data augmentation method based on the function in the Tsaug library (Wen & Keyes, 2019).
TimeGAN (Yoon et al., 2019).
@@
We also compare with the TimeGAN data augmentor, which combines supervised and adversarial objective optimization. Specifically, through a learned embedding space, the network is guided to adhere to the dynamics of the training data during sampling.
ADA (Schneider et al., 2024).

### v3

# ReAugment: Model Zoo-Guided RL for Few-Shot Time Series Augmentation and Forecasting

@@ 1 Introduction
Time series forecasting is a critical task with diverse applications, but it faces significant challenges due to the limited availability of high-quality training data. This challenge is further amplified in time-evolving domains with inherent non-stationarity and becomes even more pronounced in few-shot learning scenarios, where the scarcity of data affects the performance of forecasting models. While recent methods have focused on developing deep-learning architectures to capture long-term trends, clinical patterns, and multivariate relationships [33, 23, 21], in this work, we explore a learning-based data augmentation method that can be seamlessly integrated with existing forecasting models.
@@
Effective data augmentation requires generating high-quality, diverse training samples. However, in practice, existing forecasting models typically rely on fixed-form data augmentation techniques [30, 5], which lack data-dependent adaptability and may introduce unexpected noise. Previous learning-based augmentation methods typically involve contrastive learning [8] or Mixup combinations [26] to generate new data sequences. However, these methods are NOT task-oriented, as the data generation process is not guided by the performance of the downstream forecasting models. In contrast, we argue that enabling a closed-loop augmentation-forecasting process is essential, which requires aligning the training objectives of augmentation models with the resulting forecasting performance.
@@
To achieve this, we present ReAugment, which uses reinforcement learning (RL) to enable closed-loop data augmentation and enhance the generalizability of few-shot forecasting models. We aim to answer three critical questions: (1) Which parts of the training set should be augmented? (2) How should the augmentation be performed? (3) What advantages does RL offer compared to existing approaches? Accordingly, ReAugment is designed to dynamically identify a training subset of overfit-prone samples that would benefit most from augmentation and automatically searches for optimal augmentation policies tailored to these samples.
@@
Specifically, ReAugment initially leverages a forecasting model zoo, consisting of diverse instances of the same network architecture trained via cross-validation, to identify overfit-prone training samples in need of augmentation. An interesting finding is that training forecasting models exclusively with data samples that present low prediction variance across the model zoo often leads to better performance (detailed in Section 3). This insight leads us to select the data points with high prediction variance across the model zoo as the root of the subsequent augmentation process.
@@
Based on the identified overfit-prone samples, ReAugment learns an augmentation network in the form of a variational masked autoencoder (VMAE), in which the latent space is used as the action space in the RL framework. The RL-based augmentation policies are guided by a reward function derived from backtesting the model-zoo errors, measured by applying the generated data to the forecasting model zoo. The reward function balances both the diversity and quality of the augmented data: It encourages the new data points to fall into regions where the forecasting models tend to overfit, while maintaining alignment with the original data distribution. We employ the REINFORCE algorithm [32] to enable optimization of the VMAE under a non-differentiable reward function.
@@
In summary, we provide a pilot study on leveraging RL for time series augmentation, which demonstrates substantial improvements across a wide range of base models and datasets, especially in few-shot learning scenarios. Key technical contributions of this work are as follows:
Finding overfit-prone data as augmentation anchor points: We propose a novel method for identifying overfit-prone samples by leveraging cross-validation errors from a forecasting model zoo to pinpoint anchor points with potential value for data augmentation.
Augmentation model: We introduce a novel network architecture based on a variational masked autoencoder for sequential data, where the latent space is used as the action space of RL.
A pilot study of RL for time series augmentation: We leverage RL to incorporate a non-differentiable objective into the augmentation process. At the core of this approach is a reward function that uses prediction responses from the model zoo to guide data generation toward regions where forecasting models are likely to overfit.
@@ 2 Related Work
Compared with CNN-based or RNN-based forecasting models [28, 29, 4, 25], recent Transformer-based methods [17, 33, 40, 20, 41, 39, 38, 2, 35] have shown superior performance across a wide range of time series forecasting tasks. Most previous work has focused on developing more effective network architectures to capture long-term trends, clinical patterns, and multivariate relationships within the data. For instance, Autoformer [33] improves upon the Transformer by employing a deep decomposition architecture that progressively separates trend and seasonal components throughout the forecasting process. PatchTST [23] vectorizes time series data into patches of specified size, which are then encoded through a Transformer, with the model producing forecasts of the desired length via an appropriate prediction head. LSTF-Linear [38] simplifies complex time series forecasting problems and outperforms many Transformers by using a set of remarkably simple one-layer linear models. iTransformer [21] modifies the architecture by adopting components with inverted dimensions, demonstrating superior performance on multivariate time series data. Unlike these methods, we introduce a data augmentation method that can be broadly combined with existing forecasting models.
@@
In many real-world applications, obtaining sufficient data can be challenging, especially in scenarios where data is scarce, noisy, or difficult to collect [9, 34, 14, 37]. Recent literature has introduced large foundation models specifically designed for time series forecasting [7, 15, 1, 10, 19, 22, 24], often evaluating these models under zero-shot domain generalization settings. However, in preliminary experiments, we found that as the distribution gap between training and testing data increases (e.g., when data comes from different domains), the generalization performance of these models significantly decreases.
@@
Iglesias et al. [13] have presented a taxonomy of augmentation techniques. Simple augmentation methods involve time, frequency, and magnitude domain transformation techniques such as slicing [3], frequency warping [6], and jittering [11]. The second category is the learning-based methods, including those based on contrastive learning [8], and Mixup combinations [26]. Additionally, advanced generative models have been employed to generate realistic time series data, including the GAN-based [36, 18], VAE-based [27, 16], and Diffusion-based [12] methods. The generated samples can be used for further training of the forecasting models. In contrast to existing approaches, we present a pilot study on using RL for time series data augmentation.
@@ 3.1 Finding Overfit-Prone Data with a Model Zoo
The first challenge we need to address is identifying which parts of the training set would benefit the most, so they can be used as anchor points in the augmentation process. A key assumption in time series forecasting, particularly in few-shot learning scenarios, is that forecasting models tend to overfit certain regions of the training data. Different training samples can have varying impacts on training quality. Intuitively, we aim to generate new data with distributions around the original data points that have higher probabilities of overfitting. To achieve this, we measure the cross-validation errors from a batch of forecasting models to pinpoint the overfit-prone samples.
To construct the forecasting model zoo, we divide the training set into parts and perform -fold cross-validation. In this way, we obtain sets of model parameters for the same network architecture, e.g., iTransformer [21], which we denote as . For each data point , we evaluate the prediction errors of the other models that were not trained on this subset, along with the training error of the model trained on this subset. We then calculate the variance of the MSE across the models, denoted as , and sort the “model-zoo variance” of all data points.
The model-zoo variance measures the inconsistency in predictions made by different models trained on slightly different subsets, and can therefore serve as an indicator of the overfit-prone samples. When a model overfits, its predictions on certain data points are highly sensitive to the specific subset it was trained on, leading to a high variability of the performance across the models. As a result, data points with high model-zoo variance are more likely to be in regions where the models struggle to generalize, making them overfit-prone and thus ideal candidates for data augmentation to improve generalization. In practice, we split the training set into two subsets based on the top and bottom values of the model-zoo variance, and refer to the top subset as the overfit-prone samples.
@@ 3.2 Preliminary Findings on Overfit-Prone Data
To understand the impact of the overfit-prone samples on the training process, we conduct the following experiments. We train forecasting models separately using data from Group A (the subset with the top model-zoo variance) and Group B (the subset with the bottom model-zoo variance). We evaluate the models on the same test set. Given its superior average performance, we select iTransformer as the preferred model for these experiments. The experiments are conducted on classic real-world multivariate benchmarks, including ETT, Weather, and Electricity. The forecasting results are shown in Figure 2, where a lower MAE indicates more accurate predictions. Our experiments demonstrate that the choice of training sets significantly impacts final performance. Predictions trained in Group B consistently outperform those trained in Group A across all benchmarks, with substantial margins. This finding suggests that the overfit-prone samples are more likely to negatively affect the training quality of forecasting models. Since data augmentation is an effective strategy to mitigate overfitting, we propose generating new data around the distributions of the overfit-prone samples, allowing the forecasting model to learn more generalizable patterns from these data points.
@@ 4.2 Variational Masked Autoencoder as the RL Actor
We design a Variational Masked Autoencoder (VMAE), as illustrated in Figure 3(left), which takes masked time series data with absolute timestamp in the whole dataset as input and outputs the complete corresponding data. The entire architecture contains four modules, including (i) the prior network, which learns the prior distribution of based on the masked data, (ii) the posterior module, which learns the posterior distribution of based on the original data, (iii) the data encoder, which extracts significant features from with , (iv) the decoder, which generates data from the encoding features and the latent variables. These modules are parametrized by and :
@@
In general, the VMAE design can be integrated into any encoder-decoder-based time series forecasting architecture. In this work, we specifically adopt the encoder and decoder (implemented as a linear projector) from iTransformer [21] for feature extraction and decoding.
4.3 Variance-Guided REINFORCE Using Model Zoo Predictions
Motivation:
Few-shot time series forecasting is particularly prone to distribution imbalance and shifts, leading to sparse regions that increase overfitting risk. Thus, merely replicating the original data distribution in Stage B is insufficient. To address this, we propose an RL-based framework to expand the distribution around overfit-prone samples, aiming to better cover test patterns and avoid trivial solutions. The augmented data should (i) increase diversity near overfit-prone regions, and (ii) remain close to the original distribution to avoid noise. Inspired by our findings, we achieve this by augmenting samples that show diverse predictions across the model zoo. The key insight is to effectively use the potentially overfit-prone data to strengthen generalization.
Intuitively, the variance of prediction errors across the model zoo reflects the degree of overfitting at a given data point. The model zoo consists of network instances of the same network architecture with pretrained parameters , constructed via -fold cross-validation. Each data sample is used as test data for one model and as training data for the other models. We estimate overfitting at each training sequence by computing the variance of prediction errors across the models that were trained on it. Similarly, for the augmented data , we compute the model-zoo variance by
where represents the prediction result from the -th forecasting model. However, a practical challenge is that the model-zoo variance is non-differentiable, preventing direct gradient-based optimization of the augmentor. To address this, we finetune the neural augmentor using the REINFORCE algorithm, treating the latent space from the VMAE prior module as the action space. Specifically, we optimize the prior module as a policy network while keeping other components fixed, aiming to maximize the following reward function:
Based on this reward function, the policy network (i.e., the prior network in VMAE) is optimized as follows via gradient ascent, where is the learning rate: The reward function encourages augmented samples to exhibit more pronounced predictive variance across the model zoo, while also constraining the differences between the augmented and original data. In the reward function, a scaled-sigmoid function is employed to minimize the likelihood of rewards clustering around or controlled by the hyperparameter . This approach ensures that, despite potential order-of-magnitude differences in reconstruction error and backtest variance within the model zoo, the reward function can learn effectively from these variations.
@@ 5 Experiments
5.1 Experimental Setups
Following previous work [33, 21, 23, 38], we thoroughly evaluate the proposed ReAugment on five publicly available real-world datasets: ETT (including subsets), Traffic, Electricity, Weather, and Exchange. Please refer to Suppl. B for more details on these datasets. We evaluate ReAugment in two setups: the few-shot setup and the standard setup, which provides full access to the original training set of the above datasets.
Few-shot setup: we simulate scenarios with limited training data to assess the model’s ability to handle data scarcity. We reduce the training set size to either or of the full dataset, depending on the dataset characteristics. The few-shot training data corresponds to the earliest portion of the time series, ensuring a significant distribution shift from the test set. The validation and test sets remain consistent with those used in prior studies.
Standard setup: We follow the configuration from previous work, allowing access to the full training set while keeping the validation and test sets unchanged.
We employ a fixed lookback length of time steps across all datasets and report the multivariate sequence prediction results with prediction lengths of time steps. We provide additional implementation details in Suppl. C, including information on hyperparameters and sensitivity analysis.
Forecasting model backbones.
We primarily use iTransformer [21] for the forecasting model, as it has demonstrated strong performance in standard time series forecasting tasks. We also conduct experiments with another Transformer-based method, PatchTST [23], and the linear model DLinear [38]. Unless otherwise specified, we use a model zoo consisting of cross-validation models.
| Training Data | iTransformer | PatchTST | DLinear | Average | ||||
@@ 5.1 Experimental Setups
Compared methods.
We compare ReAugment with the following augmentation methods:
@@
Gaussian augmentation: Inspired by prior literature [13], we use traditional data augmentation methods that apply simple transformations, such as adding Gaussian noise to the raw data. This approach enhances the diversity of the training data by adjusting the controllable variances and means of the added Gaussian noise.
@@
Convolve: We employ another traditional data augmentation method based on the function in the Tsaug library [31].
@@
TimeGAN [36]: TimeGAN data augmentor combines supervised and adversarial objective optimization. Specifically, through a learned embedding space, the network is guided to adhere to the dynamics of the training data during sampling.